Jim wants to buy a 5-pound mixture of chocolates. If one kind sells for $1.80 per pound and another for $2.30 per pound, how much of each kind can he buy for $10? Here's what I have so far:180x+230(5-x)=1000,but I don't think it's right.
Answer by ptaylor(2198)

Jim wants to buy a 5-pound mixture of chocolates. If one kind sells for $1.80 per pound and another for $2.30 per pound, how much of each kind can he buy for $10? Here's what I have so far:180x+230(5-x)=1000,but I don't think it's right.


I think that you are right on target!!!!!!!! GOOD WORK!!!!!
Let x=the number of pounds of the $1.80 chocolates
Then 5-x=the number of pounds of the $2.30 chocolates
Cost of the $1.80 chocolates =$1.80x
Cost of the $2.30 chocolates =$2.30(5-x)
And we are told that the cost of the $1.80 chocolates plus the cost of the $2.30 chocolates equals $10.00. So our equation to solve is:
(Now we'll start dealing in pennies to avoid confusion----just as you did)

180x+230(5-x)=1000
180x+1150-230x=1000 Collect the x's and subtract 1150 from both sides:
-50x=-150
x=3 lbs of the $1.80 chocolates
5-x=5-3=2 lbs of the $2.30 chocolates
ck
3(180)+2(230)=1000
540+460=1000
1000=1000

Hope this helps----ptaylor